OpenStudy (anonymous):

13√12 x √75 multiply

OpenStudy (anonymous):

Are u multiplying?

OpenStudy (anonymous):

\[(13 \sqrt(12)) x \sqrt(75)\] IS THIS WHAT YOU MEAN?

OpenStudy (anonymous):

I mean what I typed.

OpenStudy (anonymous):

well i cant really tell if its a "x" variable or a multiplication sign??

OpenStudy (anonymous):

multiplying

OpenStudy (anonymous):

\[\sqrt{12\times 75}=\sqrt{2^2\times 3\times 3\times 5^2}\] \[=\sqrt{2^2\times 3^2\times 5^2}=2\times 3\times 5\]

OpenStudy (anonymous):

I still dont quite understand it.

OpenStudy (anonymous):

lets make it easier what is \(12\times 75\) ?

OpenStudy (anonymous):

900

OpenStudy (anonymous):

No it wasn't.. It was 90. thanks anyway(: